# Janus Skyrmion: Interfacial Quasiparticle with Two-Faced Helicity

### v1

@@ Abstract
Janus particles are functional particles with at least two surfaces showing asymmetric properties. We show at the interface between two magnetic regions with different antisymmetric exchange interactions, a new species of topological quasiparticles can emerge, in which different helicity structures can coexist. We name such an interfacial quasiparticle a “Janus skyrmion”, in analogy to the Janus particle. As the Janus skyrmion shows helicity asymmetry, its size could vary with both the in-plane and out-of-plane magnetic fields. A vertical spin current could drive the Janus skyrmion into one-dimensional motion along the interface without showing the skyrmion Hall effect, at a speed which depends on both the spin polarization angle and current density. Thermal fluctuations could also lead to one-dimensional random walk of a Brownian Janus skyrmion. This work uncovers unique dynamics of interfacial quasiparticles with exotic helicity structures, which may be realized in interface-engineered magnetic layers.
Introduction.—In chiral magnetic materials, antisymmetric exchange interactions, that is, Dzyaloshinskii-Moriya (DM) interactions [1, 2], could give rise to localized spin textures with nontrivial topology [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], which are usually treated as quasiparticles as they have well-defined shape, nanoscale size, and good mobility [11]. These magnetic quasiparticles could be dynamically manipulated by external drives, such as spin currents [13, 14, 15, 16, 17, 18], and are considered to be next-generation information carriers for both conventional and quantum devices [19, 20, 21, 22, 23, 24, 25, 26]. Although different magnetic quasiparticles may carry identical integer topological charge [8, 12], which is described by $Q=\frac{1}{4\pi}\int\boldsymbol{m}\cdot(\frac{\partial\boldsymbol{m}}{\partial x}\times\frac{\partial\boldsymbol{m}}{\partial y})dxdy$ with $\boldsymbol{m}$ being the reduced magnetization, their static and dynamic properties depend on their own internal structures [8, 12].
The most important internal degree of freedom of particle-like topological spin textures is their helicity [8, 12, 27, 28]. For example, among the magnetic quasiparticle family, the skyrmion is an exemplary member that can be found in ultra-thin magnetic layers and low-dimensional magnets [29, 30, 31, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], of which the helicity is mainly determined by the type of DM interactions [8, 16]. It shows Néel-type helicity in magnetic layers with interface-induced DM interactions [13, 14, 15], and shows Bloch-type helicity in magnets with bulk DM interactions (Fig. 1[29, 30, 31]. For either case, the helicity structure of a skyrmion is centrosymmetric with respect to the skyrmion center. The existence of skyrmions with helicity asymmetry as well as their static and dynamic properties have remained unexplored.
@@
The concept of Janus particles could be shifted to magnetic quasiparticles, which have at least two different or asymmetric internal physical structures. For example, skyrmions with asymmetric helicity structures may be treated as Janus quasiparticles in magnetic materials. Most recently, experimentalists have demonstrated the possibility to engineer the sign of DM interactions in magnetic layers [40, 41], which is an effective way to modify the internal structures of skyrmions and will be the key to realize Janus skyrmions.
@@
In this work, we show that one may artificially construct a Janus skyrmion with unique helicity asymmetry at the engineered interface between two magnetic regions with different types and signs of DM interactions. We explore the dynamics of such a Janus skyrmion with $|Q|=1$ driven by magnetic field, spin current, or thermal fluctuations. We find that the Janus skyrmion can be viewed as a hybridization of a Néel-type and Bloch-type skyrmions, but its mobility is limited by the one-dimensional interface, leading to exotic properties that cannot be found in Néel-type and Bloch-type skyrmions.
@@
Figure 1: Janus skyrmion quasiparticle with two-faced asymmetric helicity at the interface. (a) A basic Janus particle has two distinct surfaces, that is, the left side A and right side B, which show different or asymmetric properties. (b) The statue of the two-faced god Janus Bifrons, the ancient Roman god of beginnings and transitions, who is able to look at opposite directions. The image is a cropped version from wikipedia.org, licensed under the Creative Commons Attribution-Share Alike 3.0 Unported License (https://creativecommons.org/licenses/by-sa/3.0). (c) Top view of various relaxed Néel-type, Bloch-type, and Janus skyrmions with Q=1Q=-1. Skyrmions are fully relaxed at the center of a square magnetic layer with the same or two different types of DM interactions. The length, width, and thickness of the monolayer are equal to $40$ nm, $40$ nm, and 11 nm, respectively. The cell size is $1\times 1\times 1$ nm3. No external magnetic field is applied. For Néel-type and Bloch-type skyrmions, the DM interaction is uniform and of the same type in the model. For Janus skyrmions, the value of Néel-type DM interaction in the left region (x=020x=0-20 nm) is fixed at $D_{\text{L}}=\pm 3.7$ mJ m-2, and the strength of Bloch-type DM interaction in the right region (x=2040x=20-40 nm) is set to $|D_{\text{R}}|=|D_{\text{L}}|$. The type and sign of $D_{\text{L}}$ and $D_{\text{R}}$ are indicated beyond each skyrmion configuration. Other default parameters are: $\alpha=0.3$, $M_{\text{S}}=580$ kA m-1, $A=15$ pJ m-1, and $K=0.8$ MJ m-3. The color scale represents the reduced out-of-plane ($m_{z}$) or in-plane ($m_{x}$) magnetization component. The magnetization configurations are indicated by black arrows. (d) A basic Janus skyrmion has two distinct helicity structures. For example, the left half is with Néel-type helicity, and the right half is with Bloch-type helicity. The shape of the Janus skyrmion is thus like a heart, which is in stark contrast to a normal skyrmion with centrosymmetric helicity. (e) Topological mapping of the spin configurations of a two-dimensional Janus skyrmion with Q=1Q=-1 [(d)] onto a three-dimensional 22-sphere.
@@
From Janus particle to Janus skyrmion.—We first study the static structure of a Janus skyrmion. A Janus particle has two faces with different properties [Fig. 1(a)], therefore, we focus on a Janus skyrmion with two different helicity structures, as illustrated in Figs. 1(c) and 1(d). The left half of the Janus skyrmion is of Néel-type helicity, while the right half is of Bloch-type helicity. Such a helicity structure is not centrosymmetric with respect to the skyrmion center, which is different to skyrmions with intermediate helicity between Néel-type and Bloch-type ones [42, 44, 45, 46, 43]. Consequently, the Janus skyrmion with helicity asymmetry has a unique heart-like shape but carries a topological charge carried of $|Q|=1$.
@@
We assume that a Janus skyrmion with Q=1Q=-1 and two different helicity structures is formed at the interface between two magnetic regions with different types of DM interactions. In our simulation [Fig. 1(c)], we place and relax a skyrmion at the center of a magnetic layer. The size of the magnetic layer is $40\times 40\times 1$ nm3, which includes $1,600$ spins. The spin dynamics is governed by the Landau-Lifshitz-Gilbert (LLG) equation $\partial_{t}\boldsymbol{m}=-\gamma_{0}\boldsymbol{m}\times\boldsymbol{h}_{\text{eff}}+\alpha(\boldsymbol{m}\times\partial_{t}\boldsymbol{m})$, which is solved by using the OOMMF simulator [47]. In the LLG equation, the absolute gyromagnetic ratio $\gamma_{0}=2.211\times 10^{5}$ m A-1 s-1 and the damping parameter $\alpha=0.3$. The effective field $\boldsymbol{h}_{\text{eff}}=-\frac{1}{\mu_{0}M_{\text{S}}}\cdot\frac{\delta\varepsilon}{\delta\boldsymbol{m}}$, where $\mu_{0}$ and $\varepsilon$ denote the vacuum permeability constant and average energy density, respectively. The average energy density includes the contributions from the ferromagnetic exchange interaction, DM exchange interactions (Néel-type and Bloch-type), perpendicular magnetic anisotropy (PMA), demagnetization, and applied magnetic field, given as $\varepsilon=A\left(\nabla\boldsymbol{m}\right)^{2}+D_{\text{N}}\left[m_{z}\left(\boldsymbol{m}\cdot\nabla\right)-\left(\nabla\cdot\boldsymbol{m}\right)m_{z}\right]+D_{\text{B}}\left[\boldsymbol{m}\cdot\left(\nabla\times\boldsymbol{m}\right)\right]-K(\boldsymbol{n}\cdot\boldsymbol{m})^{2}-\frac{M_{\text{S}}}{2}(\boldsymbol{m}\cdot\boldsymbol{B}_{\text{d}})-M_{\text{S}}(\boldsymbol{m}\cdot\boldsymbol{B}_{\text{a}})$. Note that $D_{\text{N}}$ and $D_{\text{B}}$ stand for the Néel-type and Bloch-type DM interaction parameters, respectively. $\boldsymbol{B}_{\text{d}}$ is the demagnetization field, $\boldsymbol{B}_{\text{a}}$ is the applied magnetic field, and $\boldsymbol{n}$ is the unit surface normal vector. The default magnetic parameters are [16, 17]: the saturation magnetization $M_{\text{S}}=580$ kA m-1, the exchange constant $A=15$ pJ m-1, and PMA constant $K=0.8$ MJ m-3.
@@
As shown in Fig. 1(c), if we consider uniform Néel-type or Bloch-type DM interaction in the magnetic layer (either positive or negative), the relaxed skyrmion is of Néel-type or Bloch-type, where the helicity structure is centrosymmetric with respect to the skyrmion center. However, if we assume a Néel-type DM interaction in the left half of the magnetic layer ($x\in[0~\text{nm},20~\text{nm}]$), and a Bloch-type DM interaction in the right half of the magnetic layer ($x\in[20~\text{nm},40~\text{nm}]$), the relaxed skyrmion at the interface between the two different magnetic regions shows hybrid and asymmetric helicity structures, including both Néel-type and Bloch-type spin configurations. Such a Janus skyrmion is stable at the $D_{\text{N}}$-$D_{\text{B}}$ interface and its topological charge is Q=1Q=-1 [Fig. 1(e)]. Its asymmetric internal spin configurations can be modified by changing the sign of DM interaction in each region [Fig. 1(c)]. In Supplementary Fig. 1 [48], we also show that the size and shape of the Janus skyrmion depend on the strength of the DM interaction in each region. The Janus skyrmion may be deformed when the DM interactions in the left and right regions are not equal to each other.
@@
In-plane and out-of-plane fields.—In Fig. 2, we apply an external magnetic field $\boldsymbol{B}_{\text{d}}$ to a Janus skyrmion, and observe its response to the field. The Janus skyrmion is initially relaxed in a magnetic layer of $100\times 100\times 1$ nm3, which includes $10,000$ spins. The DM interaction in the left half of the magnetic layer is of Néel-type, and it is of Bloch-type in the right half. The relaxed Janus skyrmion at zero magnetic field shows asymmetric helicity, as given in Fig. 1(d).
@@
In theory, a normal Néel-type or Bloch-type skyrmion [Fig. 1(c)] cannot be manipulated by a moderate in-plane magnetic field, as the magnetic field forces acting on the in-plane spins of the skyrmion will cancel with each other due to the centrosymmetry of the helicity structure. Only a strong in-plane magnetic field may deform the skyrmion into a noncircular shape by dragging perpendicularly magnetized spins into the $x$-$y$ plane [49]. However, we find that the size of the Janus skyrmion with helicity asymmetry can be adjusted by an in-plane magnetic field. In Fig. 2, an external magnetic field applied in the x-x or y-y direction could increase the skyrmion size, while a field applied in the +x+x or +y+y direction could shrink the skyrmion. This is a unique feature that cannot be found in normal circular skyrmions with centrosymmetric helicity structures. Indeed, similar to the case of a normal skyrmion, an out-of-plane magnetic field applied in the $\pm z$ direction can modify the size of the Janus skyrmion, which is an expected result.
@@
Figure 3: Current-induced dynamics of a Janus skyrmion at the interface ($\theta$ dependence). (a) The velocity of the Janus skyrmion along the interface (i.e., $\pm y$ directions) as functions of the spin polarization angle ($\theta$). Symbols are simulation results, and curves show theoretical solutions. (b) The $\theta$-dependent speed of the Janus skyrmion moving along the interface. The driving current density $j$ is fixed at 11 or 33 MA cm-2. Only simulation results are shown. Here, the Janus skyrmion is stabilized at the interface between Néel-type (left) and Bloch-type (right) DM interactions. $D_{\text{L}}=D_{\text{R}}=3.5$ mJ m-2. The model size is $100\times 100\times 1$ nm3. Other parameters are given in Fig. 1 caption.
@@
Figure 4: Current-induced dynamics of a Janus skyrmion at the interface ($j$ dependence). (a) Left: The velocity of the Janus skyrmion along the interface (i.e., $\pm y$ directions) as a function of the driving current density ($j$). Right: The reduced out-of-plane magnetization of the system obtained during the stable motion ($m_{z}^{s}$) as a function of $j$. The spin polarization angle is fixed at $\theta=270$ degrees (i.e., $\boldsymbol{p}=-\hat{y}$). (b) The time-dependent $m_{z}$ during the motion of a Janus skyrmion driven by different current densities (j=110j=1-10 MA cm-2). See Fig. 3 caption for other parameters.
Figure 5: One-dimensional Brownian random walk of a Janus skyrmion at the interface. (a) Time-dependent position of the Janus skyrmion in the $x$ dimension. (b) Time-dependent position of the Janus skyrmion in the $y$ dimension. (c) Time-dependent $m_{z}$ of the model. Here, the Brownian dynamics of the Janus skyrmion is simulated at three different temperatures ($T=10$, $50$, and $100$ K). See Fig. 3 caption for other parameters.

### v2

@@ Abstract
Janus particles are functional particles with at least two surfaces showing asymmetric properties. We show at the interface between two magnetic regions with different antisymmetric exchange interactions, a new species of two-dimensional topological quasiparticles can emerge, in which different helicity structures can coexist. We name such an interfacial quasiparticle a “Janus skyrmion”, in analogy to the Janus particle. As the Janus skyrmion shows helicity asymmetry, its size could vary with both the in-plane and out-of-plane magnetic fields. A vertical spin current could drive the Janus skyrmion into one-dimensional motion along the interface without showing the skyrmion Hall effect, at a speed which depends on both the in-plane spin polarization direction and current density. Thermal fluctuations could also lead to one-dimensional random walk of a Brownian Janus skyrmion. This work uncovers unique dynamics intrinsic to interfacial quasiparticles with exotic helicity, which may be realized in interface-engineered magnetic layers.
I Introduction
In chiral magnetic materials, antisymmetric exchange interactions, that is, the Dzyaloshinskii-Moriya (DM) interactions [1, 2], could give rise to localized spin textures with nontrivial topology [3, 4, 5, 6, 7, 8, 9, 10, 13, 11, 12], which are usually treated as quasiparticles as they have well-defined shape, nanoscale size, and good mobility [13]. These magnetic quasiparticles could be dynamically manipulated by external drives, such as spin currents [13, 14, 16, 17, 18, 19, 20, 21, 15, 22, 23], and are considered to be next-generation information carriers for both conventional and quantum devices [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Although different magnetic quasiparticles may carry identical integer topological charge [8, 11], which is described by $Q=\frac{1}{4\pi}\int\boldsymbol{m}\cdot(\frac{\partial\boldsymbol{m}}{\partial x}\times\frac{\partial\boldsymbol{m}}{\partial y})dxdy$ with $\boldsymbol{m}$ being the reduced magnetization, their static and dynamic properties depend on their internal structures [8, 11].
In principle, the size, shape, and internal structures of topological spin textures can be modified by external means [12]. An important internal degree of freedom of particle-like topological spin textures is their helicity [8, 11, 40, 41, 39, 42, 35, 38, 36, 37]. For example, among the magnetic quasiparticle family, the skyrmion is an exemplary member that can be found in ultra-thin magnetic layers and low-dimensional magnets [43, 44, 45, 3, 4, 5, 6, 7, 8, 9, 10, 13, 11, 12], of which the helicity is mainly determined by the type of DM interactions [8, 19]. It shows Néel-type helicity in magnetic layers with interface-induced DM interactions [16, 17, 18], and shows Bloch-type helicity in magnets with bulk DM interactions (Fig. 1[43, 44, 45]. For either case, the helicity structure of a two-dimensional (2D) skyrmion is centrosymmetric with respect to the skyrmion center. The 2D skyrmions with helicity asymmetry as well as their static and dynamic properties remain insufficiently explored.
@@ I Introduction
The concept of Janus particles could be shifted to magnetic quasiparticles, which have at least two different or asymmetric internal physical structures. For example, 2D skyrmions with asymmetric helicity structures may be treated as Janus quasiparticles in magnetic materials. Most recently, experimentalists have demonstrated the possibility to engineer the sign of the DM interaction [54, 55], which is an effective way to induce helicity asymmetry in 2D skyrmions and will be the key to realize Janus skyrmions.
@@
In this work, we show that one may artificially construct a Janus skyrmion with unique helicity asymmetry at the engineered interface between two magnetic regions with different types and signs of DM interactions. We explore the dynamics of such a Janus skyrmion driven by magnetic fields, spin currents, or thermal fluctuations. We find that the Janus skyrmion can be viewed as a hybridization of Néel-type and Bloch-type skyrmions. However, its mobility is limited by the one-dimensional (1D) interface, leading to exotic properties that cannot be found in conventional Néel-type and Bloch-type skyrmions.
@@
Figure 1: Janus skyrmion quasiparticle with two-faced asymmetric helicity at the interface. (a) A basic Janus particle has two distinct surfaces, that is, the left side A and right side B, which show different or asymmetric properties. (b) The statue of the two-faced god Janus Bifrons, the ancient Roman god of beginnings and transitions, who is able to look at opposite directions. The image is a cropped version from wikipedia.org, licensed under the Creative Commons Attribution-Share Alike 3.0 Unported License. (c) Top view of various relaxed Néel-type, Bloch-type, and Janus skyrmions with Q=1Q=-1. Skyrmions are fully relaxed at the center of a square magnetic layer with the same or two different types of DM interactions. The length, width, and thickness of the monolayer are equal to $40$ nm, $40$ nm, and 11 nm, respectively. The cell size is $1\times 1\times 1$ nm3. No external magnetic field is applied. For Néel-type and Bloch-type skyrmions, the DM interaction is uniform and of the same type in the model. For Janus skyrmions, the value of Néel-type DM interaction in the left region (x=020x=0-20 nm) is fixed at $D_{\text{L}}=\pm 3.7$ mJ m-2, and the strength of Bloch-type DM interaction in the right region (x=2040x=20-40 nm) is set to $|D_{\text{R}}|=|D_{\text{L}}|$. The type and sign of $D_{\text{L}}$ and $D_{\text{R}}$ are indicated beyond each skyrmion configuration. A minimum value of $|D_{\text{R}}|=|D_{\text{L}}|=3.1$ mJ m-2 is required to stabilize the Janus skyrmion. The color scale represents the reduced out-of-plane ($m_{z}$) or in-plane ($m_{x}$) spin component. The spin configurations are indicated by black arrows. (d) A basic Janus skyrmion has two distinct helicity structures. For example, the left half is with Néel-type helicity, and the right half is with Bloch-type helicity. The shape of the Janus skyrmion is thus like a heart, which is in stark contrast to a conventional skyrmion with centrosymmetric helicity. (e) Topological mapping of the spin configurations of a 2D Janus skyrmion with Q=1Q=-1 [(d)] onto a three-dimensional (3D) 22-sphere.
@@
II Results and Discussion
II.1 From Janus particle to Janus skyrmion
@@
We first study the static structure of a Janus skyrmion. A Janus particle has two faces with different properties [Fig. 1(a)], therefore, we focus on a Janus skyrmion with two different helicity structures, as illustrated in Figs. 1(c) and 1(d). The left half of the Janus skyrmion is of Néel-type helicity, while the right half is of Bloch-type helicity. Such a hybrid helicity structure is not centrosymmetric with respect to the skyrmion center, which is different to skyrmions with intermediate helicity between Néel-type and Bloch-type ones [56, 57, 58, 59, 42]. Consequently, the Janus skyrmion with helicity asymmetry carries a topological charge of $|Q|=1$ but has a unique heart-like shape. Previously, heart-like skyrmions are also found in chiral skyrmions with chiral kinks [37] as well as in modified multilayer magnets [60].
@@
We assume that a Janus skyrmion with Q=1Q=-1 and two different helicity structures is formed at the interface between two magnetic regions with different types of DM interactions. In our simulation [Fig. 1(c)], we place and relax a skyrmion at the center of a magnetic layer. The size of the magnetic layer is $40\times 40\times 1$ nm3, which includes $1,600$ spins. The spin dynamics is governed by the Landau-Lifshitz-Gilbert (LLG) equation $\partial_{t}\boldsymbol{m}=-\gamma_{0}\boldsymbol{m}\times\boldsymbol{h}_{\text{eff}}+\alpha(\boldsymbol{m}\times\partial_{t}\boldsymbol{m})$, which is solved by using the OOMMF simulator [61]. In the LLG equation, the absolute gyromagnetic ratio $\gamma_{0}=2.211\times 10^{5}$ m A-1 s-1 and the damping parameter $\alpha=0.3$ (for dynamics simulation). The effective field $\boldsymbol{h}_{\text{eff}}=-\frac{1}{\mu_{0}M_{\text{S}}}\cdot\frac{\delta\varepsilon}{\delta\boldsymbol{m}}$, where $\mu_{0}$ and $\varepsilon$ denote the vacuum permeability constant and average energy density, respectively. The average energy density includes the contributions from the ferromagnetic exchange interaction, DM interactions (Néel-type and Bloch-type), perpendicular magnetic anisotropy (PMA), demagnetization, and applied magnetic field, given as $\varepsilon=A\left(\nabla\boldsymbol{m}\right)^{2}+D_{\text{N}}\left[m_{z}\left(\boldsymbol{m}\cdot\nabla\right)-\left(\nabla\cdot\boldsymbol{m}\right)m_{z}\right]+D_{\text{B}}\left[\boldsymbol{m}\cdot\left(\nabla\times\boldsymbol{m}\right)\right]-K(\boldsymbol{n}\cdot\boldsymbol{m})^{2}-\frac{M_{\text{S}}}{2}(\boldsymbol{m}\cdot\boldsymbol{B}_{\text{d}})-M_{\text{S}}(\boldsymbol{m}\cdot\boldsymbol{B}_{\text{a}})$. Note that $D_{\text{N}}$ and $D_{\text{B}}$ stand for the Néel-type (i.e., interfacial-like) and Bloch-type (i.e., bulk-like) DM interaction parameters, respectively. $\boldsymbol{B}_{\text{d}}$ is the demagnetization field, $\boldsymbol{B}_{\text{a}}$ is the applied magnetic field, and $\boldsymbol{n}$ is the unit surface normal vector. The default magnetic parameters are [19, 20, 15, 62, 63, 64]: the saturation magnetization $M_{\text{S}}=580$ kA m-1, the exchange constant $A=15$ pJ m-1, and PMA constant $K=0.8$ MJ m-3. These magnetic parameters are based on an as-deposited ultra-thin magnetic film (e.g., cobalt) on a heavy-metal layer (e.g., platinum) grown by sputtering deposition [15], which can effectively stabilize nanoscale skyrmions. For modeling purposes, we consider different types of DM interactions in the left and right regions [Fig. 1(d)]. The interfacial-like DM interaction in the left half of the model is of the $C_{nv}$ symmetry group [65] and the corresponding DM interaction vectors ($\vec{d}_{ij}$) are given in a counterclockwise configuration (i.e., $\vec{d}_{ij}=\vec{u}_{ij}\times\hat{z}$ with $\vec{u}_{ij}$ being the unit vector between site $i$ and $j$, and $\hat{z}$ being the surface normal). In the right half of the model, the bulk-like DM interaction is of the $D_{n}$ symmetry group [55] and the corresponding DM vectors are given in a radical configuration (i.e., $\vec{d}_{ij}=\vec{u}_{ij}$). In principle, the configuration of DM interaction vectors can be modified by controlling symmetry, spin-orbit coupling, and local atomic environment at the interface between the magnetic layer and heavy-metal substrate. In experiments, these modifications can be realized by applying an electric field on a local region, inducing strain locally, or changing the interface composition or stacking order. Particularly, to change the DM interaction vector configuration from interfacial-like to bulk-like, one can apply an electric field [42, 35, 38, 36, 32, 33] only to the right half of the sample in a partially voltage-gated device structure. A recent experimental report on voltage-controlled DM interactions has demonstrated the possibility to deform Néel-type skyrmions into Bloch-type ones by applying an electric field to control the DM interaction vectors [42]. The Janus skyrmions could be realized at the boundary between voltage-controlled and uncontrolled regions. Also, a previous study suggested that one can construct a heart-shaped Janus skyrmion by utilizing the edge effect of the oxide layer upon a skyrmion-hosting magnetic layer [60].
@@
As shown in Fig. 1(c), if we consider uniform Néel-type or Bloch-type DM interaction in the magnetic layer (either positive or negative), the relaxed skyrmion is of Néel-type or Bloch-type, where the helicity structure is centrosymmetric with respect to the skyrmion center. However, if we assume a Néel-type DM interaction in the left half of the magnetic layer ($x\in[0~\text{nm},20~\text{nm}]$), and a Bloch-type DM interaction in the right half of the magnetic layer ($x\in[20~\text{nm},40~\text{nm}]$), the relaxed skyrmion at the interface between the two different magnetic regions shows hybrid and asymmetric helicity structures, including both Néel-type and Bloch-type spin configurations. Such a Janus skyrmion is stable at the $D_{\text{N}}$-$D_{\text{B}}$ interface and its topological charge is Q=1Q=-1 [Fig. 1(e)]. Its asymmetric internal spin configurations can be modified by changing the signs of $D_{\text{N}}$ and $D_{\text{B}}$ [Fig. 1(c)]. Note that the Janus skyrmion stabilized by positive $D_{\text{N}}$ and $D_{\text{B}}$ could be transformed to that stabilized by negative $D_{\text{N}}$ and $D_{\text{B}}$ through the symmetry operation ($m_{x}$, $m_{y}$, $m_{z}$) $\rightarrow$ ($m_{x}$, $m_{y}$, $m_{z}$). In Supplementary Fig. 1 [66], we show that the Janus skyrmion also exhibits non-uniform distributions of both topological charge density and total energy density, which are in sharp contrast to that of conventional Néel-type and Bloch-type skyrmions. Moreover, the size and shape of the Janus skyrmion depend on the DM interaction strength in each region. It may be deformed when the DM interactions in the left and right regions are not identical (Supplementary Fig. 2 [66]).
II.2 In-plane and out-of-plane magnetic fields
@@
In Fig. 2, we apply an external magnetic field $\boldsymbol{B}_{\text{a}}$ to a Janus skyrmion, and observe its response to the field. The Janus skyrmion is initially relaxed in a magnetic layer of $100\times 100\times 1$ nm3, which includes $10,000$ spins. The DM interaction in the left half of the magnetic layer is of Néel-type, and it is of Bloch-type in the right half. The relaxed Janus skyrmion at zero magnetic field shows asymmetric helicity, as given in Fig. 1(d).
@@
In theory, a conventional Néel-type or Bloch-type skyrmion [Fig. 1(c)] cannot be manipulated by an in-plane magnetic field, as the magnetic field forces acting on the in-plane spins of the skyrmion will cancel with each other due to the centrosymmetry of the helicity structure. Only a strong in-plane magnetic field may deform the skyrmion into a noncircular shape by dragging perpendicularly magnetized spins into the $x$-$y$ plane [67]. However, we find that the size of the Janus skyrmion with helicity asymmetry can be adjusted by an in-plane magnetic field. In Fig. 2, an external magnetic field applied in the x-x or y-y direction could increase the skyrmion size, while a field applied in the +x+x or +y+y direction could shrink the skyrmion. This is a unique feature that cannot be found in conventional circular skyrmions with centrosymmetric helicity structures. Indeed, similar to the case of a conventional skyrmion, an out-of-plane magnetic field applied in the $\pm z$ direction can modify the size of the Janus skyrmion, which is an expected result. We note that the light stripe along the interface between the left and right regions, visible in Fig. 2, is a result of the difference between the DM interactions in the left and right regions. The DM interactions lead to micromagnetic boundary conditions that tilt boundary spins at the interface [65]. Namely, the interfacial-like DM interaction ensures that the boundary spins rotate in a plane containing the boundary surface normal, while the bulk-like DM interaction ensures that the boundary spins rotate in a plane parallel to the boundary surface [65]. Consequently, due to the uniform symmetric ferromagnetic exchange coupling in the model, the spins exhibit a smooth variation across the interface in the perpendicularly magnetized layer, while a tiny in-plane spin component arises from the different tilt directions of the spins adjacent to the interface.
@@
Figure 3: Current-induced dynamics of a Janus skyrmion at the interface ($\theta$ dependence). (a) The velocity of the Janus skyrmion along the interface (i.e., $\pm y$ directions) as functions of the spin polarization angle ($\theta$). Symbols are simulation results, and curves show theoretical solutions. (b) The $\theta$-dependent speed of the Janus skyrmion moving along the interface. The current density $j$ is fixed at 11 or 33 MA cm-2. Only simulation results are shown. Here, the Janus skyrmion is stabilized at the interface between Néel-type (left) and Bloch-type (right) DM interactions. $D_{\text{L}}=D_{\text{R}}=3.5$ mJ m-2. The model size is $100\times 100\times 1$ nm3. Other parameters are given in Fig. 1 caption.
@@
Figure 4: Current-induced dynamics of a Janus skyrmion at the interface ($j$ dependence). (a) Left: The velocity of the Janus skyrmion along the interface (i.e., $\pm y$ directions) as a function of the current density ($j$). Right: The reduced out-of-plane magnetization of the system obtained during the stable motion ($m_{z}^{s}$) as a function of $j$. The spin polarization angle is fixed at $\theta=270$ degrees (i.e., $\boldsymbol{p}=-\hat{y}$). (b) The time-dependent $m_{z}$ during the motion of a Janus skyrmion driven by different current densities (j=110j=1-10 MA cm-2). See Fig. 3 caption for other parameters.
II.3 Current-induced dynamics